The measure of the angles of a triangle are three consecutive even integers. Find the measure of each angle.
Found 2 solutions by edjones, Alan3354:
Answer by edjones(8007)

Let the angles be x, x+2, x+4
x+x+2+x+4=180
3x+6=180
3x=174
x=58
.
Ed

Answer by Alan3354(69443)

The 3 angles add to 180.
180/3 = 60, the middle number.
--> 58, 60, 62